# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ Appendix B Iron core and mantle densities including volatiles
The Water Sequestration model presented here simplifies the treatment of the interior structure to determine the radius of the iron core and magma ocean by assuming a water layer on top of the rocky material instead of consistently mixing water with mantle and core. In this section, we aim at comparing the radii of the planets obtained in this way with a more consistent model based on Dorn et al. (2017).
This model focuses on Earth-like rocky interiors with steam atmospheres. Water can exist in the core, mantle, or surface, depending on the thermal state of the planet. The core is composed of iron (Fe) with light elements such as hydrogen (H) and oxygen (O). For solid Fe, we use the equations of state for hexagonal close-packed iron from Hakim et al. (2018), and for liquid iron and its alloys, we follow Luo et al. (2024) who use a Mie-Grüneisen equation of state with parameters derived from ab-initio calculations. Luo et al. (2024) state that their approach is valid in the super-Earth to sub-Neptune regime, that is for high pressures above 50 GPa. The core is assumed to be adiabatic, but there is a temperature jump at the core-mantle boundary (CMB) due to residual heat from core formation, based on Stixrude (2014).
The mantle consists of three primary components: MgO, SiO2, and FeO, assuming an Earth-like composition. For solid mantle properties, we employ the Perple_X thermodynamic model Connolly (2009) and the database of Stixrude & Lithgow-Bertelloni (2022). For pressures above 125 GPa, stable minerals are defined a priori using equations of state from various sources (Fischer et al. 2011; Faik et al. 2018; Hemley et al. 1992; Musella et al. 2019; Coppari et al. 2021). The liquid mantle’s density is computed as an ideal mixture of Mg2SiO4, SiO2, and FeO, using Mg2SiO4 instead of MgO due to the recent data for forsterite being updated for high-pressure regimes.
The mantle is considered fully adiabatic, with water present only in the melts, while solid mantle is assumed to be dry. Water decreases the mantle’s density following Bajgain et al. (2015), and this reduction is nearly pressure and temperature independent for small water fractions. The mantle melting curve is calculated for pure MgSiO3, and the addition of water (Katz et al. 2003) and iron (Dorn et al. 2018) lowers the melting temperature. Water in the core also reduces the melting temperature, as modeled by Luo et al. (2024). The water which is not dissolved in the mantle (see Sect. 2.2.1) is modeled as being in steam or supercritical phases using the AQUA equation of state (Haldemann et al. 2020), with an isothermal profile below 0.1 bar that transitions to an adiabatic profile.
Hydrated silicates are not considered motivated by a minor impact on the density (Shah et al. 2021). Instead, water is treated using the EOS of Haldemann et al. (2020) and the additive volume law to combine it with rock. In this way, interior structures and planetary core and mantle radii can be obtained comparable to the simplified model adopted in this work.

### v2

@@ Appendix B Iron core and mantle densities including volatiles
The Water Sequestration model presented here simplifies the treatment of the interior structure to determine the radius of the iron core and magma ocean by assuming a water layer on top of the rocky material instead of consistently mixing water with mantle and core. In this section, we aim at comparing the radii of the planets obtained in this way with a more consistent model based on Dorn2017.
This model focuses on Earth-like rocky interiors with steam atmospheres. Water can exist in the core, mantle, or surface, depending on the thermal state of the planet. The core is composed of iron (Fe) with light elements such as hydrogen (H) and oxygen (O). For solid Fe, we use the equations of state for hexagonal close-packed iron from Hakim2018, and for liquid iron and its alloys, we follow Luo2024 who use a Mie-Grüneisen equation of state with parameters derived from ab-initio calculations. Luo2024 state that their approach is valid in the super-Earth to sub-Neptune regime, that is for high pressures above 50 GPa. The core is assumed to be adiabatic, but there is a temperature jump at the core-mantle boundary (CMB) due to residual heat from core formation, based on Stixrude2014.
The mantle consists of three primary components: MgO, SiO2, and FeO, assuming an Earth-like composition. For solid mantle properties, we employ the Perple_X thermodynamic model Connolly2009 and the database of Stixrude2022. For pressures above 125 GPa, stable minerals are defined a priori using equations of state from various sources (Fischer2011; Faik2018; Hemley1992; Musella2019; Coppari2021). The liquid mantle’s density is computed as an ideal mixture of Mg2SiO4, SiO2, and FeO, using Mg2SiO4 instead of MgO due to the recent data for forsterite being updated for high-pressure regimes.
The mantle is considered fully adiabatic, with water present only in the melts, while solid mantle is assumed to be dry. Water decreases the mantle’s density following Bajgain2015, and this reduction is nearly pressure and temperature independent for small water fractions. The mantle melting curve is calculated for pure MgSiO3, and the addition of water (Katz2003) and iron (Dorn2018) lowers the melting temperature. Water in the core also reduces the melting temperature, as modeled by Luo2024. The water which is not dissolved in the mantle (see Sect. 2.2.1) is modeled as being in steam or supercritical phases using the AQUA equation of state (Haldemann2020), with an isothermal profile below 0.1 bar that transitions to an adiabatic profile.
Hydrated silicates are not considered motivated by a minor impact on the density (Shah2021). Instead, water is treated using the EOS of Haldemann2020 and the additive volume law to combine it with rock. In this way, interior structures and planetary core and mantle radii can be obtained comparable to the simplified model adopted in this work.